# Worth Their Weight: Randomized and Regularized Block Kaczmarz Algorithms without Preprocessing

### v1

@@ 5.1 Numerical Demonstration
To test our algorithms for calculating natural gradient directions, we train a small neural network to learn a simple function on the unit interval. We take three snapshots from the training process and form the least-squares problem (31) for each. The methods TA-mSGD, TA-RBK-U, and TA-ReBlocK-U are then compared on these problems with results in Figure 3. The computations are performed on an A100 GPU to simulate a deep learning setting, and progress is measured with wall-clock time on the horizontal axis. TA-ReBlocK-U performs best in every case, achieving an accuracy nearly an order of magnitude better than TA-RBK-U in the first snapshot and a full order of magnitude better than TA-mSGD in the second snapshot. Further details on these experiments can be found in Section F.2. For additional context, we examine the singular values of each problem and confirm that they exhibit exponential decay in the top part of the spectrum; see Section F.2.2.
Our results, though preliminary due to their small scale, suggest that ReBlocK is a promising method for calculating natural gradient directions. This provides justification for the Kaczmarz-inspired SPRING algorithm (Goldshlager et al., 2024) and suggests a family of related methods to be explored by future works. Open questions include how many ReBlock iterations to run between each update of $\theta$, how to incorporate averaging, and how to tune $\lambda$.
@@ 6 Conclusions
In this work, we have explored the problem of solving large-scale linear least-squares problems without preprocessing the data matrix $A$. Our results suggest that ReBlocK may be a more effective algorithm than mSGD for inconsistent problems that exhibit rapid singular value decay. More broadly, our work highlights the value of incorporating regularization as a path towards broadening the applicability of block Kaczmarz methods, and our analysis suggests that a Monte Carlo perspective can be useful in elucidating the behavior of randomized block row-access methods in general. Finally, our work provides motivation and suggests new directions for least-squares based natural gradient optimizers.
@@ Acknowledgements
G.G. acknowledges support from the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0023112. This effort was supported by the SciAI Center, and funded by the Office of Naval Research (ONR), under Grant Number N00014-23-1-2729 (J.H., L.L.). L.L. is a Simons Investigator in Mathematics. This research used the Savio computational cluster resource provided by the Berkeley Research Computing program at the University of California, Berkeley. We thank Ethan Epperly, Robert Webber, and Ruizhe Zhang for their thoughtful discussions and feedback.
Disclaimer
@@ Appendix A Helpful Lemmas
$\mathbb{E}\left[x_{s}-x^{(\rho)}|x_{r}\right]=(I-\overline{P})^{s-r}(x_{r}-x^{(\rho)}).$ (32)
Proof of Lemma A.1.
@@
$\mathbb{E}\left[x_{t+1}-x^{(\rho)}|x_{t}\right]=(I-\overline{P})(x_{t}-x^{(\rho)})+A^{\top}\overline{W}r^{(\rho)}=(I-\overline{P})(x_{t}-x^{(\rho)}).$ (34)
Using the law of total expectation and the linearity of expectation, iterating this result for $t=r,\ldots,s-1$ yields the lemma. ∎
@@
$x^{\top}(I-\overline{P})^{s}x\leq(1-\alpha)^{s}\left\lVert x\right\rVert^{2}$ (36)
@@
$\left\lVert(I-\overline{P})^{s}x\right\rVert\leq(1-\alpha)^{s}\left\lVert x\right\rVert.$ (37)
Proof.
@@
$\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{t}B+V$ (38)
@@
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{1}{\alpha(T-T_{b})}V.$ (39)
Proof.
We follow closely the Proof of Theorems 1.2 and 1.3 of (Epperly et al., 2024). Decompose the expected mean square error as
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}=\frac{1}{(T-T_{b})^{2}}\sum_{r,s=T_{b}+1}^{T}\mathbb{E}\left[(x_{r}-x^{(\rho)})^{\top}(x_{s}-x^{(\rho)})\right].$ (40)
@@
where the third line uses Lemmas A.2 and A.3 and the assumptions $\operatorname{range}{\overline{P}}=\operatorname{range}(A^{\top})$ and $\overline{P}\preceq I$.
@@
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{V}{(T-T_{b})^{2}}\sum_{r,s=T_{b}+1}^{T}(1-\alpha)^{|s-r|}.$ (41)
@@
$\sum_{r,s=T_{b}+1}^{T}(1-\alpha)^{|s-r|}\leq 2\sum_{r=T_{b}+1}^{T}\sum_{s=0}^{\infty}(1-\alpha)^{s}=\frac{T-T_{b}}{\alpha}$ (42)
@@
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{V}{\alpha(T-T_{b})}.$ (43)
@@
$\left\lVert r^{(\rho)}\right\rVert^{2}\leq\kappa(\overline{W})\left\lVert r\right\rVert^{2}.$ (44)
Proof.
@@ Appendix B Proofs of RBK Convergence Theorems
In this section we provide the proofs of Theorems 3.1, 3.2 and 3.3. We first prove the following lemma, which can be viewed as a more general version of Theorem 1.2 of (Needell & Tropp, 2014):
@@
$\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{t}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}+\frac{1}{\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}.$ (45)
Proof.
@@
Iterating from $s=0$ to $s=t-1$ and utilizing $\sum_{s=0}^{t-1}(1-\alpha)^{s}<\sum_{s=0}^{\infty}(1-\alpha)^{s}=\nicefrac{{1}}{{\alpha}}$ yields the desired result. ∎
Proof of Theorem 3.1.
The result follows from the appropriate application of Lemmas A.1, A.4 and B.1. To apply these lemmas it is first required to verify $\overline{P}\preceq I$ and $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$. For the first result, it suffices to observe that $P(S)$ is the orthogonal projector onto the row space of $A_{S}$ and thus $P(S)\preceq I$ for all $S$.
@@
$x^{\top}\overline{P}x=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[x^{\top}A_{S}^{\top}(A_{S}A_{S}^{\top})^{-1}A_{S}x\right]\geq\frac{k}{m}\cdot\frac{(a_{i}^{\top}x)^{2}}{\left\lVert a_{i}\right\rVert^{2}}>0,$ (47)
@@
$\mathbb{E}\left[x_{T}\right]-x^{(\rho)}=(I-\overline{P})^{T}(x_{0}-x^{(\rho)}).$ (48)
@@
$\left\lVert\mathbb{E}\left[x_{T}\right]-x^{(\rho)}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{(\rho)}\right\rVert$ (49)
using Lemmas A.2 and A.3 and the proven properties of $\overline{P}$.
Furthermore, Lemma B.1 holds and provides the conditions for Lemma A.4 with $B=\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}$, $V=\frac{1}{\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}$. The application of Lemma A.4 then directly implies the convergence result for the tail-averaged RBK-U algorithm. ∎
Proof of Theorem 3.2.
@@
$y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a_{i}^{\top}\;b_{i}]\sim\mathcal{N}(0,Q)}\left[(a_{i}^{\top}x-b_{i})^{2}\right].$ (50)
Furthermore, the solution $x^{*}$ is the corresponding (minimal-norm) empirical risk minimizer. It is a standard result in asymptotic statistics that as the number of samples $m$ grows, the empirical minimizer $x^{*}$ converges to the true solution $y$ (Lehmann & Casella, 2006; Van der Vaart, 2000).
@@
$y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a_{i}^{\top}\;b_{i}]\sim\mathcal{N}(0,Q)}\left[(a_{i}^{\top}x-b_{i})^{2}\right]=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\|L^{\top}x-d\|^{2}.$ (52)
@@
We note that a similar independence lemma has also been established in Lemma 3.14 of (Rebrova & Needell, 2021). Furthermore, taking the expectation on both sides of (53), we have
@@
$\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$ (54)
In addition, the variance term $\left\lVert A_{S_{t}}^{\top}r_{S_{t}}\right\rVert^{2}$ in Theorem 3.1 takes the form
$\left\lVert(Z^{\top}L^{\top})^{+}Z^{\top}\tilde{r}\right\rVert^{2}.$ (55)
@@
$\left\lVert(Z^{\top}L^{\top})^{+}Z^{\top}\tilde{r}\right\rVert^{2}=\left\lVert(Z_{1}^{\top}L^{\top})^{+}Z_{2}^{\top}\tilde{r}\right\rVert^{2}\leq\sigma_{\rm min}(Z_{1}^{\top}L^{\top})^{-2}\left\lVert Z_{2}^{\top}\tilde{r}\right\rVert^{2}.$ (56)
We can bound the expectation of the smallest singular value of $Z_{1}^{\top}L^{\top}$ using a similar technique to the proof of Lemma 22 in (Dereziński & Rebrova, 2024). To start, let $L^{\top}=W\Sigma Y^{\top}$ be the SVD of $L^{\top}$. Then
$\sigma_{\rm min}(Z_{1}^{\top}L^{\top})^{2}=\sigma_{\rm min}(Z_{1}^{\top}L^{\top}LZ_{1})=\sigma_{\rm min}(Z_{1}^{\top}W\Sigma^{2}W^{\top}Z_{1}).$ (57)
@@
Since the columns of $W_{2k}$ are orthonormal the random matrix $Z_{1}^{\top}W_{2k}$ can be redistributed as a single Gaussian random matrix $G_{2k}$ of size $k\times(2k)$ and with standard normal entries. By Lemma 3.16 of (Rebrova & Needell, 2021), we thus have for $k\geq 6$,
$\mathbb{E}\left[\sigma_{\rm min}^{-2}(Z_{1}^{\top}L^{\top})\right]\leq\frac{20}{(\sqrt{2k}-\sqrt{k})^{2}\sigma_{2k}^{2}}\leq\frac{200}{k\sigma_{2k}^{2}}.$ (58)
@@
It remains only to bound the value of $\alpha$. Note that the entries of $Z_{t}\in\mathbb{R}^{d\times m}$ are independently drawn from the standard Gaussian distribution.
This property enables us to leverage existing results from (Dereziński & Rebrova, 2024) concerning the spectrum of the matrix $\overline{P}$. Specifically, $P_{t}=(Z_{t}L^{\top})^{+}Z_{t}L^{\top}$, which has a similar formula to the matrix $P$ considered in Equation (6) of (Dereziński & Rebrova, 2024). Then, applying Theorem 3.1 of (Dereziński & Rebrova, 2024) gives the result stated in Theorem 3.2. ∎
Proof of Corollary 3.3..
By Corollary 3.4 of (Dereziński & Rebrova, 2024), if $L$ has the polynomial spectral decay of order $\beta>1$, i.e., $\sigma_{i}^{2}\leq i^{-\beta}\sigma_{1}^{2}$ for all $i$, the dependence of $C_{n,k}$ on $n$ and $k$ in Theorem 3.2 can be eliminated when $k\leq n/2$. Furthermore, there is a constant $C=C(\beta)$ such that for any $k\leq n/2$, the linear convergence rate satisfies
@@
Regarding the convergence rate of mSGD, let us consider solving a fixed number of samples, say $m$, of $\{a_{i}^{\top},b_{i}\}$. It is demonstrated in (Needell et al., 2014) that, with a minibatch size of 1 and importance sampling, the corresponding convergence parameter can be at most $\alpha^{\rm SGD}\leq 1/\kappa_{\rm dem}(\frac{1}{m}A_{m}A_{m}^{\top})$ with $A_{m}=[a_{1},\dots,a_{m}]$ for convergent step sizes. When a larger minibatch size $k$ is used, (Jain et al., 2018b) shows that the learning rate can be increased at most linearly, corresponding to a rate of at most $\alpha^{\rm mSGD}\leq k/\kappa_{\rm dem}(\frac{1}{m}A_{m}A_{m}^{\top})$. It follows by the random matrix theory (Johnstone, 2001; Tao, 2012) that $\lim_{m\rightarrow\infty}\frac{1}{m}A_{m}A_{m}^{\top}=LL^{\top}$. Hence, when $m$ goes to infinity, the linear convergence rate of mSGD satisfies

### v2

@@ 5.1 Numerical Demonstration
Figure 5: Comparison of methods for calculating natural gradient directions for a small neural network. The network parameters $\theta$ are taken from three snapshots of a single training run, with one snapshot from the “pre-descent” phase before the loss begins to decrease (top left), one snapshot from the “descent” phase during which the loss decreases rapidly (top right), and one snapshot from the “post-descent” phase when the decay rate of the loss has slowed significantly (bottom). The algorithms are measured in terms of their progress towards reducing the relative residual $\tilde{r}=\left\lVert Jx-[f_{\theta}-f]\right\rVert/\left\lVert f_{\theta}-f\right\rVert$ for the least-squares problem (32), which measures how well the function-space update direction $Jx$ agrees with the function-space loss gradient $f_{\theta}-f$. The burn-in time is set to $T_{b}\approx T/100$ in each case, as indicated by the vertical dotted line.
To test our algorithms for calculating natural gradient directions, we train a neural network with about $7.5\times 10^{5}$ parameters to learn a rapidly oscillating function on the unit interval. We take three snapshots from the training process and form the least-squares problem (32) for each. The methods TA-mSGD, TA-RBK-U, and TA-ReBlocK-U are then compared on these problems with results in Fig. 5. The computations are performed on an A100 GPU to simulate a deep learning setting, and progress is measured with wall-clock time on the horizontal axis. TA-ReBlocK-U performs best in every case, achieving an accuracy over an order of magnitude better the other methods in both the second and third snapshots.
In Fig. 6, we investigate the effect of the parameter $\lambda$ on the performance of ReBlocK. We observe that choosing $\lambda=1e-1$ already provides a significant speed-up relative to mSGD, and the speed-up is further improved by reducing $\lambda$ down to $31e-$ 3. However, choosing $\lambda=1e-4$ results in unstable behavior similar to the RBK algorithm. These results suggest that the performance of ReBlocK is optimized by choosing $\lambda$ as small as possible without introducing significant instabilities. Next, in Fig. 7 we explore the effect of the batch size $k$ on the performance of the algorithms. We observe that ReBlocK performs well across all batch sizes, although RBK is competitive for the largest batch size of $k=200$. Finally, in Fig. 8, we report the number of iterations per second for mSGD, RBK, and ReBlocK for the experiments of this section. We find that the ReBlocK iterations are over twenty-five times faster than the RBK iterations and only about $6\%$ slower than the mSGD iterations, which is a significant reason that TA-ReBlocK-U outperforms TA-RBK-U in practice. Further details on these experiments can be found in F.2.
Figure 6: Performance of TA-ReBlocK-U on the problem from the top right panel of Fig. 5 with various values of $\lambda$.
Figure 7: Performance of algorithms for different values of the block size $k$. The problem is the same as in the top right panel of Fig. 5 and the values of $k$ are $k=25$ (top left), $k=50$ (top right), $k=100$ (bottom left), and $k=200$ (bottom right). The top right panel here is identical to the top right panel of Fig. 3, and is reproduced here only for comparison with the other quadrants.
Figure 8: Iteration speed of each algorithm for the natural gradient experiments with $k=50$.
Our results, though preliminary due to their synthetic nature, suggest that ReBlocK is a promising method for calculating natural gradient directions. This provides justification for the Kaczmarz-inspired SPRING algorithm [19] and suggests a family of related methods to be explored by future works. Open questions include how many ReBlock iterations to run between each update of $\theta$, whether to incorporate averaging, and how to tune $\lambda$.
@@ 6 Conclusions
In this work, we have explored the problem of solving large-scale linear least-squares problems without preprocessing the data matrix $A$. Our results suggest that ReBlocK is a more effective algorithm than either RBK or mSGD for inconsistent problems that exhibit rapid singular value decay. More broadly, our work highlights the value of incorporating regularization as a path towards broadening the applicability of block Kaczmarz methods, and our analysis suggests that a Monte Carlo perspective can be useful in elucidating the behavior of randomized block row-access methods in general. Finally, our work provides motivation and suggests new directions for least-squares based natural gradient optimizers.
@@ Acknowledgements
G.G. acknowledges support from the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0023112. This effort was supported by the SciAI Center, and funded by the Office of Naval Research (ONR), under Grant Number N00014-23-1-2729 (J.H., L.L.). L.L. is a Simons Investigator in Mathematics. This research used the Savio computational cluster resource provided by the Berkeley Research Computing program at the University of California, Berkeley. We thank Ethan Epperly, Robert Webber, Ruizhe Zhang, and anonymous reviewers for their thoughtful discussions and feedback.
@@ Appendix A Helpful Lemmas
(33) $\mathbb{E}\left[x_{s}-x^{(\rho)}|x_{r}\right]=(I-\overline{P})^{s-r}(x_{r}-x^{(\rho)}).$
Proof A.2 (Proof of Lemma A.1).
@@
(35) $\mathbb{E}\left[x_{t+1}-x^{(\rho)}|x_{t}\right]=(I-\overline{P})(x_{t}-x^{(\rho)})+A^{\top}\overline{W}r^{(\rho)}=(I-\overline{P})(x_{t}-x^{(\rho)}).$
Using the law of total expectation and the linearity of expectation, iterating this result for $t=r,\ldots,s-1$ yields the lemma.
@@
(37) $x^{\top}(I-\overline{P})^{s}x\leq(1-\alpha)^{s}\left\lVert x\right\rVert^{2}$
@@
(38) $\left\lVert(I-\overline{P})^{s}x\right\rVert\leq(1-\alpha)^{s}\left\lVert x\right\rVert.$
Proof A.6.
@@
(39) $\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{t}B+V$
@@
(40) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{1}{\alpha(T-T_{b})}V.$
Proof A.8.
We follow closely the Proof of Theorems 1.2 and 1.3 of [18]. Decompose the expected mean square error as
(41) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}=\frac{1}{(T-T_{b})^{2}}\sum_{r,s=T_{b}+1}^{T}\mathbb{E}\left[(x_{r}-x^{(\rho)})^{\top}(x_{s}-x^{(\rho)})\right].$
@@
where the third line uses Lemmas A.3 and A.5 and the assumptions $\operatorname{range}{\overline{P}}=\operatorname{range}(A^{\top})$ and $\overline{P}\preceq I$.
@@
(42) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{V}{(T-T_{b})^{2}}\sum_{r,s=T_{b}+1}^{T}(1-\alpha)^{|s-r|}.$
@@
(43) $\sum_{r,s=T_{b}+1}^{T}(1-\alpha)^{|s-r|}\leq 2\sum_{r=T_{b}+1}^{T}\sum_{s=0}^{\infty}(1-\alpha)^{s}=\frac{T-T_{b}}{\alpha}$
@@
(44) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{V}{\alpha(T-T_{b})}.$
@@
(45) $\left\lVert r^{(\rho)}\right\rVert^{2}\leq\kappa(\overline{W})\left\lVert r\right\rVert^{2}.$
Proof A.10.
@@ Appendix B Proofs of RBK Convergence Theorems
In this section we provide the proofs of Theorems 3.1, 3.2, and 3.3. We first prove the following lemma, which can be viewed as a more general version of Theorem 1.2 of [31]:
@@
(46) $\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{t}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}+\frac{1}{\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}.$
Proof B.2.
@@
Iterating from $s=0$ to $s=t-1$ and utilizing $\sum_{s=0}^{t-1}(1-\alpha)^{s}<\sum_{s=0}^{\infty}(1-\alpha)^{s}=\nicefrac{{1}}{{\alpha}}$ yields the desired result.
Proof B.3 (Proof of Theorem 3.1).
The result follows from the appropriate application of Lemmas A.1, A.7, and B.1. To apply these lemmas it is first required to verify $\overline{P}\preceq I$ and $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$. For the first result, it suffices to observe that $P(S)$ is the orthogonal projector onto the row space of $A_{S}$ and thus $P(S)\preceq I$ for all $S$.
@@
(48) $x^{\top}\overline{P}x=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[x^{\top}A_{S}^{\top}(A_{S}A_{S}^{\top})^{-1}A_{S}x\right]\geq\frac{k}{m}\cdot\frac{(a_{i}^{\top}x)^{2}}{\left\lVert a_{i}\right\rVert^{2}}>0,$
@@
(49) $\mathbb{E}\left[x_{T}\right]-x^{(\rho)}=(I-\overline{P})^{T}(x_{0}-x^{(\rho)}).$
@@
(50) $\left\lVert\mathbb{E}\left[x_{T}\right]-x^{(\rho)}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{(\rho)}\right\rVert$
using Lemmas A.3 and A.5 and the proven properties of $\overline{P}$.
Furthermore, Lemma B.1 holds and provides the conditions for Lemma A.7 with $B=\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}$, $V=\frac{1}{\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}$. The application of Lemma A.7 then directly implies the convergence result for the tail-averaged RBK-U algorithm.
Proof B.4 (Proof of Theorem 3.2).
@@
(51) $y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a_{i}^{\top}\;b_{i}]\sim\mathcal{N}(0,Q)}\left[(a_{i}^{\top}x-b_{i})^{2}\right].$
Furthermore, the solution $x^{*}$ is the corresponding (minimal-norm) empirical risk minimizer. It is a standard result in asymptotic statistics that as the number of samples $m$ grows, the empirical minimizer $x^{*}$ converges to the true solution $y$ [25, 44].
@@
(53) $y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a_{i}^{\top}\;b_{i}]\sim\mathcal{N}(0,Q)}\left[(a_{i}^{\top}x-b_{i})^{2}\right]=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\|L^{\top}x-d\|^{2}.$
@@
We note that a similar independence lemma has also been established in Lemma 3.14 of [37]. Furthermore, taking the expectation on both sides of (54), we have
@@
(55) $\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$
Now, consider the bound on the value of $\alpha$. Note that the entries of $Z_{t}\in\mathbb{R}^{d\times m}$ are independently drawn from the standard Gaussian distribution. This property enables us to leverage existing results from [14] concerning the spectrum of the matrix $\overline{P}$. Specifically, $P_{t}=(Z_{t}L^{\top})^{+}Z_{t}L^{\top}$, which has a similar formula to the matrix $P$ considered in Equation (6) of [14]. Then, applying Theorem 3.1 of [14] gives the result stated in Theorem 3.2.
It remains to show the bound on the variance. Note that the variance term $\left\lVert A_{S_{t}}^{\top}r_{S_{t}}\right\rVert^{2}$ in Theorem 3.1 takes the form
(56) $\left\lVert(Z^{\top}L^{\top})^{+}Z^{\top}\tilde{r}\right\rVert^{2}.$
@@
(57) $\left\lVert(Z^{\top}L^{\top})^{+}Z^{\top}\tilde{r}\right\rVert^{2}=\left\lVert(Z_{1}^{\top}L^{\top})^{+}Z_{2}^{\top}\tilde{r}\right\rVert^{2}\leq\sigma_{\rm min}(Z_{1}^{\top}L^{\top})^{-2}\left\lVert Z_{2}^{\top}\tilde{r}\right\rVert^{2}.$
We can bound the expectation of the smallest singular value of $Z_{1}^{\top}L^{\top}$ using a similar technique to the proof of Lemma 22 in [14]. To start, let $L^{\top}=W\Sigma Y^{\top}$ be the SVD of $L^{\top}$. Then
(58) $\sigma_{\rm min}(Z_{1}^{\top}L^{\top})^{2}=\sigma_{\rm min}(Z_{1}^{\top}L^{\top}LZ_{1})=\sigma_{\rm min}(Z_{1}^{\top}W\Sigma^{2}W^{\top}Z_{1}).$
@@
Since the columns of $W_{2k}$ are orthonormal the random matrix $Z_{1}^{\top}W_{2k}$ can be redistributed as a single Gaussian random matrix $G_{2k}$ of size $k\times(2k)$ and with standard normal entries. By Lemma 3.16 of [37], we thus have for $k\geq 6$,
(59) $\mathbb{E}\left[\sigma_{\rm min}^{-2}(Z_{1}^{\top}L^{\top})\right]\leq\frac{20}{(\sqrt{2k}-\sqrt{k})^{2}\sigma_{2k}^{2}}\leq\frac{200}{k\sigma_{2k}^{2}}.$
@@
Proof B.5 (Proof of Corollary 3.3.).
By Corollary 3.4 of [14], if $L$ has the polynomial spectral decay of order $\beta>1$, i.e., $\sigma_{i}^{2}\leq i^{-\beta}\sigma_{1}^{2}$ for all $i$, the dependence of $C_{n,k}$ on $n$ and $k$ in Theorem 3.2 can be eliminated when $k\leq n/2$. Furthermore, there is a constant $C=C(\beta)$ such that for any $k\leq n/2$, the linear convergence rate satisfies
@@
Regarding the convergence rate of mSGD, let us consider solving a fixed number of samples, say $m$, of $\{a_{i}^{\top},b_{i}\}$. It is demonstrated in [32] that, with a minibatch size of 1 and importance sampling, the corresponding convergence parameter can be at most $\alpha^{\rm SGD}\leq 1/\kappa_{\rm dem}(\frac{1}{m}A_{m}A_{m}^{\top})$ with $A_{m}=[a_{1},\dots,a_{m}]$ for convergent step sizes. When a larger minibatch size $k$ is used, [23] shows that the learning rate can be increased at most linearly, corresponding to a rate of at most $\alpha^{\rm mSGD}\leq k/\kappa_{\rm dem}(\frac{1}{m}A_{m}A_{m}^{\top})$. It follows by the random matrix theory [24, 43] that $\lim_{m\rightarrow\infty}\frac{1}{m}A_{m}A_{m}^{\top}=LL^{\top}$. Hence, when $m$ goes to infinity, the linear convergence rate of mSGD satisfies